jhanavi wants to make a rectangular park for children and others to play . the area of the park remains unchanged if its length is increased by 7m and breadth is decreased by 3m unit. its area remains uneffected if the length is decreased by 7m and breadth is increased by 5m.find the dimentions of the park why did jhanvi decide to make a park.
Answer by richwmiller(17219)

(x+7)*(y-3)=x*y,
(x-7)*(y+5)=x*y
x = 28, y = 15
check
x*y=420
35*12=420
21*20=420
He has money to give away and enjoys math problems!
Park dog millionaire